. How long will it take the tool to fall to the ground? (Hint: use the first

equation that you were given above, 2

0 ht t h ( ) 16 =− + . For the building’s

height, use the height of the building that you estimated in Task 1.) 


1
Expert's answer
2021-10-26T14:53:00-0400

h(t)=-16t2+h0, we are given height of the building is 70ft,so at 0, h(t)=70,so we can use it h(0)=70, h(0)=-16(0)2+h0 "\therefore" 70=-16(0)2+h0 so h0=70. Now we can plug it back so we can set h(t)=0 and then we can solve for t , h(t)=-16t2+70, t2=35/8 "\therefore" t= 2.09165